OpenStudy (anonymous):

find all the zeroes of the quadratic poly. abx^2+(b^2-a)

OpenStudy (anonymous):

@apoorvk

OpenStudy (anonymous):

@Diyadiya

OpenStudy (apoorvk):

So equate this thing to zero. \[abx^2 + (b^2-a) = 0\]\[or,~ x^2 = \frac{a - b^2}{ab}\]\[or, x = \pm\sqrt{\frac{a - b^2}{ab}}\] Does this help?

OpenStudy (anonymous):

can u make it into de stndrd form,,,, ax^2+bx+c=0

OpenStudy (apoorvk):

Oh yeah, we can work that way too.. comparing coefficients with std. form \(Ax^2 + Bx + C=0\) what would A, B and C be according to you?

OpenStudy (anonymous):

i guess a=ab, b=0, c=b^2-a??

OpenStudy (apoorvk):

yes!! So, use this is in the quadratic formula: \[x = \frac{-B \pm \sqrt{B^2 - 4AC}}{2A}\]

OpenStudy (apoorvk):

Since B = 0, you will get the same thing basically. So, this quadratic formula process is similar to taking the route around the globe! o.O

OpenStudy (anonymous):

itz damn so confusing...let ma teachr suffr de rrest....helpin me teach dis ;)

OpenStudy (apoorvk):

since B=0, \(\large x = \frac{\sqrt{-4AC}}{2A} = \sqrt{\frac {-4AC}{4A^2}} = \sqrt{\frac {-C}{A}}\) Does this make it simple?

OpenStudy (anonymous):

ummm....nice try... bt ma head is kindah dumb to undrstnd dis....itz k jz leav it...